# Quantum querying based on multicontrolled Toffoli gates for causal Feynman loop configurations and directed acyclic graphs

### v1

@@ 5.3 Five eloops with ten egdes
The last topology we implement is a five-eloop case with a single edge per set $s_{i}$, Fig. 6E. The clauses to validate are composed by
@@
The output from the algorithm gives an angle $\theta=28.9^{\circ}$ with a single iteration $t=1$, without the need of fixing a qubit from the register $|{e}\rangle$. This means that the probability of search success is close to one, $\sin^{2}(\theta_{t})=0.99$ , finding all the $240$ causal states out of the $2^{10}$ total states. Note that in this case we need to test every clause in both directions. With respect of the requirements for building the quantum circuit, we need 17 qubits to construct the corresponding quantum circuit (Fig. 9), and the quantum circuit depth is equal to 42.
@@ 6 Conclusions
We have presented an improved quantum algorithm for the identification of causal configurations for multiloop Feynman diagrams, which is equivalent to querry DAG configurations of undirected graphs in graph theory. The approach presented in this paper construct the oracle operator through multicontrolled Toffoli gates and XNOT gates only, allowing to reduce the total number of qubits required significantly. Additionaly, the quantum circuit depth do not increase by adding extra edges per set, it depends exclusively on the eloop clauses. Due to the lowering on the number of qubits needed, it was possible to explore more complex multiloop topologies through the IBM Quantum platform, in particular, we explicitly implemented the three eloops with six and nine vertices, the four eloops with twelve and sixteen vertices, and the five eloops with five point interaction with ten edges, successfully identifying their causal configurations.
@@ Acknowledgements
We are very grateful to Fundación Centro Tecnológico de la Información y la Comunicación (CTIC) for granting us access to their simulator Quantum Testbed (QUTE). We thank also access to IBMQ. This work is supported by the Spanish Government (Agencia Estatal de Investigación MCIN/AEI/ 10.13039/501100011033) Grant No. PID2020-114473GB-I00, and Generalitat Valenciana Grant No. PROMETEO/2021/071 and ASFAE/2022/009 (Planes Complementarios de I+D+i, NextGenerationEU). This work is also supported by the Ministry of Economic Affairs and Digital Transformation of the Spanish Government and NextGenerationEU through the Quantum Spain project, and by CSIC Interdisciplinary Thematic Platform (PTI+) on Quantum Technologies (PTI-QTEP+). SRU acknowledges support from CONAHCyT through Project No. 320856 (Paradigmas y Controversias de la Ciencia 2022), Ciencia de Frontera 2021-2042 and Sistema Nacional de Investigadores; AERO from the Spanish Government (PRE2018-085925).

### v2

@@ 5.3 Five eloops with ten egdes
Five-eloop topology, Fig. 9E
Quantum device (qubits) Transpiled quantum depth Two-qubit gates
BC MCX BC MCX
ibm_brisbane (127) 10745 13008 3629 4377
ibm_kyiv (127) 12410 13909 3629 4377
ibm_sherbrooke (127) 12485 13281 3629 4377
Table 7: Key metrics for the transpilation of the quantum circuit corresponding to the four-eloop topology of Fig. 9E in different quantum devices for the binary-clause (BC) and multicontrolled Toffoli (MCX) algorithms.
The last eloop topology we implement has five eloops with a single edge per set $s_{i}$ which is depicted in Fig. 9E. The eloop clauses to validate are composed by
@@
The mixing angle is $\theta=28.9^{\circ}$ without the need of fixing a qubit from the register $|{e}\rangle$. This means that the querying probability for finding the 240 causal states out of the $2^{10}$ total states is close to one, $\sin^{2}(\theta_{t})=0.99$, with a single iteration, $t=1$.
In terms of quantum resources, the algorithm based on multi-controlled Toffoli gates requires seventeen qubits to build the corresponding quantum circuit (Fig. 12) with a quantum circuit depth of twenty-three versus thirty-one qubits and a quantum circuit depth of twenty-five. Note that in this case the inclusion of additional edges in the five-eloop topology would prevent the binary-clause algorithm from being implemented in the Qiskit quantum simulator, given its limited capacity of 32 qubits.
Continuing the analysis of the five-eloop topology we follow with the study of the transpiled quantum circuit depth for relevant metrics in different quantum backends, both simulated and real quantum hardware, as shown in Table 7. The transpiled quantum circuit depth is larger and requires more two-qubit gates for the multicontrolled Toffoli algorithm than for the binary-clause algorithm.
Figure 13: Quantum circuit depth (upper plot) and quantum circuit area (bottom plot) for the five-eloop topology of Fig. 9E in terms of optimisation levels for the transpilation. The solid lines represent the transpilation of the BC quantum algorithm and the dashed lines represent the transpilation for the MCX.
As mentioned in the previous section, to provide a more appropriate assessment of the complexity of the algorithm we use the quantum circuit area considering both, the transpiled quantum depth and qubit usage. In Fig. 13 we present the transpiled quantum circuit depth and the transpiled quantum circuit area as a function of the optimisation levels on different quantum hardware, noting comparable results with the four-eloop topology. It shows a similar evolution in terms of the optimisation levels, and despite having a transpile quantum depth higher for the MCX algorithm, the quantum circuit area is smaller compared to the BC algorithm.
In addition, the simulations revealed a significant difference in execution time and resource consumption. The time taken in the ideal simulation in Qiskit (noiseless) was around 165 seconds for the BC algorithm and less than 0.1 seconds for the MCX algorithm. The BC algorithm simulation fully utilises the computer processor (100% CPU load on an Intel i7 processor, running Qiskit local on a Linux operating system) and consumed $\sim 90\%$ of the available 64 GB RAM, pushing the limits of the quantum simulator. In contrast, the MCX algorithm simulation was executed with significantly fewer resources.
@@ 6 Conclusions
We have presented an efficient and systematic quantum algorithm for identifying the causal configurations for multiloop Feynman diagrams, which is equivalent to querying DAG configurations of undirected graphs in graph theory. The approach presented in this paper defines an oracle operator using only multicontrolled Toffoli gates and NOT (Pauli-$X$) gates simplifying the design and encoding of the quantum circuits. Furthermore, it allows us to significantly reduce the total number of qubits required and the theoretical quantum circuit depth with respect to a binary-clause algorithm.
We have also analysed the quantum complexity of the transpiled quantum circuits. In this case, the quantum circuit area is the most appropriate metric to assess the complexity of the quantum algorithm due to the fact that jointly evaluates the number of qubits and the transpiled quantum depth. The quantum circuit area of the proposed algorithm is smaller than that of the binary-clause algorithm, despite having a greater transpiled quantum circuit depth, thus showing a relevant improvement in quantum resource efficiency. The reduction in the number of qubits required has allowed us to explore complex multiloop topologies using the IBM Quantum platform; in particular, we have explicitly implemented three eloops with six and nine vertices, four eloops with twelve and sixteen vertices, and five eloops with a five point interaction and ten edges, successfully identifying their causal configurations.
@@ Acknowledgements
We are very grateful to Fundación Centro Tecnológico de la Información y la Comunicación (CTIC) for granting us access to their simulator Quantum Testbed (QUTE). We thank also access to IBMQ. This work is supported by the Spanish Government (Agencia Estatal de Investigación MCIN/AEI/ 10.13039/501100011033) Grant No. PID2020-114473GB-I00, No. PID2023-146220NB-I00 and No. CEX2023-001292-S, and Generalitat Valenciana Grant No. PROMETEO/2021/071 and ASFAE/2022/009 (Planes Complementarios de I+D+i, NextGenerationEU). This work is also supported by the Ministry of Economic Affairs and Digital Transformation of the Spanish Government and NextGenerationEU through the Quantum Spain project, and by CSIC Interdisciplinary Thematic Platform (PTI+) on Quantum Technologies (PTI-QTEP+). SRU acknowledges support from CONAHCyT through Project No. 320856 (Paradigmas y Controversias de la Ciencia 2022), Ciencia de Frontera 2021-2042 and Sistema Nacional de Investigadores; AERO from the Spanish Government (PRE2018-085925).